OpenStudy (anonymous):

Does anyone like Trig identities? Verify the trig identity: cos2x/1+ sin2x = cotx - 1/cotx+1 Need the work shown, I dont know how to execute the problem

OpenStudy (anonymous):

To answer your first question, no.

OpenStudy (mertsj):

\[\frac{\cot x-1}{\cot x+1}=\frac{\frac{\cos x}{\sin x}-1}{\frac{\cos x}{\sin x}+1}=\frac{\cos x-\sin x}{\cos x+\sin x}\] \[\frac{\cos x-\sin x}{\cos x+\sin x}\times \frac{\cos x+\sin x}{\cos x+\sin x}=\frac{\cos ^{2}x-\sin ^{2}x}{\cos ^{2}x+2\sin x \cos x+\sin ^{2}x}\]

OpenStudy (mertsj):

\[\frac{\cos ^2x-\sin ^2x}{\cos ^2x+2\sin x \cos x+\sin ^2x}=\frac{\cos 2x}{1+2\sin x \cos x}=\frac{\cos 2x}{1+\sin 2x}\]

OpenStudy (anonymous):

Thank you so much this really helped me figure it out!! :)

OpenStudy (mertsj):

I can help you in a little bit. I am helping someone else right now. Sorry